# DistRL: An Asynchronous Distributed Reinforcement Learning Framework for On-Device Control Agents

### v1

@@ 1 Introduction
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications Alayrac et al. (2022); Achiam et al. (2023); Zheng et al. (2024); Li et al. (2023). This versatility enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios Yang et al. (2023b); Wang et al. (2024). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
To bridge this gap, fine-tuning(type of light training) aforementioned agents on demonstrations via imitation learning Zhan & Zhang (2023); Nakano et al. (2021); Yang et al. (2023b) channels pre-trained abilities into actionable behaviors suitable for device control. However, the dynamic nature of devices and the web renders models trained on static data sub-optimal as ecosystems evolve. Such agents struggle to recover from mistakes in changing environments, limiting their practical utility. Therefore, building robust and reliable device-control agents necessitates an interactive approach that enables MLLMs to adapt and learn from their own experiences on devices and the Internet.
Moreover, MLLMs are prone to confidently generating incorrect content that deviates from human preferences Ouyang et al. (2022); Ziegler et al. (2019); Bai et al. (2022a); Stiennon et al. (2020). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from Human Feedback (RLHF), becomes essential. RLHF leverages human feedback to align model outputs with desired behaviors and preferences. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to human expectations. Details of the problem formulation as an RL-based fine-tuning approach will be discussed in Sections 2.2 and 3.
@@ 2.1 Multi-Modal On-device Control Agents
Recent advancements in pre-trained Large Language Models (LLMs) and Multimodal LLMs (MLLMs) have revolutionized on-device control agents, moving beyond early methods like behavioral cloning or reinforcement learning Osa et al. (2018); Mnih et al. (2015); Shi et al. (2017). Early agents simulated mouse clicks and typing Shi et al. (2017); Humphreys et al. (2022) but faced scalability and adaptability challenges.
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT Nakano et al. (2021) employ fine-tuned models for web browsing, while WebAgent Gur et al. (2023) generates web code using T5. AppAgent Yang et al. (2023b) and MobileAgent Wang et al. (2024) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding Driess et al. (2023); Reid et al. (2024), but without fine-tuning, they are limited by the base models Driess et al. (2023).
Some works fine-tune VLMs with demonstration data, e.g., AutoUI, CogAgent Kapoor et al. (2024); Zhan & Zhang (2023), but static data-trained models may struggle with real-world variability Jiang et al. (2023). Others use filtered imitation learning with autonomously collected data Pan et al. (2024); Lai et al. (2024), While DigiRL Bai et al. (2024) supports on-device RL fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes. Additionally, DigiRL lacks support for efficient distributed learning algorithm designs, further hindering its scalability and performance in dynamic, parallel settings. To address this, our scalable and asynchronous RL fine-tuning pipeline, DistRL, offers an efficient solution for distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences Stiennon et al. (2020); Ouyang et al. (2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations Ouyang et al. (2022); Bai et al. (2022a).
Recent advances in MLLMs Alayrac et al. (2022); Driess et al. (2023); Reed et al. (2022); Li et al. (2023); Zhang et al. (2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences Ouyang et al. (2022); Ziegler et al. (2019); Bai et al. (2022a); Stiennon et al. (2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative Bai et al. (2022b); Yu et al. (2024); Lee et al. (2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots Bai et al. (2024); Lee et al. (2023); Yu et al. (2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems Reed et al. (2022); Liang et al. (2023); Driess et al. (2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods Driess et al. (2023); Nakano et al. (2021); Reed et al. (2022); Yao et al. (2023). This approach is effective for large-scale applications like Android device control.
@@ 5.1 A-RIDE: The Backbone of DistRL
While GAE has proven highly effective in synchronous environments, its reliance on synchronized data collection makes it less suitable for asynchronous settings, such as distributed control tasks on devices. To handle asynchronous trajectory generation in device control tasks, A-RIDE leverages an enhanced Retrace($\lambda$) method for robust off-policy corrections inspired by Espeholt et al. (2018); Munos et al. (2016). Unlike traditional on-policy algorithms that require synchronous data collection, A-RIDE is designed for distributed environments with asynchronous data.
@@
$\mathcal{L}=-\mathbb{E}_{\mu}\left[\rho_{t}A(s_{t},a_{t})\log\pi(a_{t}|s_{t})\right]-\beta\mathbb{E}_{\mu}\left[H(\pi(a_{t}|s_{t}))\right]+\lambda\mathbb{E}_{\mu}\left[\mathcal{P}_{\text{invalid}}(a_{t})\right],$ (2)
where $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini Reid et al. (2024)), ensuring that only contextually appropriate actions are penalized. The hyperparameters $\beta$ and $\lambda$ are optimized through empirical studies to balance exploration and policy robustness effectively. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness.
@@ 5.2 Distributed Prioritized Experience Replay (DPER)
To improve sample efficiency, we employ Distributed Prioritized Experience Replay (DPER). For each trajectory $\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{t=0}^{H}$, we compute the priority $p(\tau)$ as: $p(\tau)=w_{1}\overline{|\delta|}+w_{2}\overline{\rho}+w_{3}\overline{H},$ where $\overline{|\delta|}$ is the average absolute temporal-difference (TD) error over the trajectory, calculated as $\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$; $\overline{\rho}$ is the average importance sampling ratio $\rho_{t}$; and $\overline{H}$ is the average policy entropy, $H_{t}=-\log\pi(a_{t}|s_{t})$, encouraging exploration by encouraging policy uncertainty, thus avoiding early convergence to suboptimal policies during training in dynamic environments. The weights $w_{1}$, $w_{2}$, and $w_{3}$ balance the contributions of each component, which is selected by grid-search. Trajectories with higher priorities are replayed more frequently, focusing learning on the most informative experiences. Priorities are periodically updated based on the latest policy, recalculating them to focus learning on the most informative experiences, ensuring continual adaptation to evolving behavior policies. Details can be found in Appendix A.4.
@@ A.4.2 Implementation Details
Policy Update Mechanism
The actor (policy network) is updated using gradients derived from the advantage estimates:
$\nabla_{\theta}\mathcal{L}=-\mathbb{E}_{\mu}\left[\rho_{t}A(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right]-\beta\nabla_{\theta}\mathbb{E}_{\mu}\left[\log\pi_{\theta}(a_{t}|s_{t})\right],$ (6)
where $\theta$ represents the parameters of the policy network.
@@ A.4.3 Hyperparameter Tuning for Distributed Prioritized Experience Replay
These ranges were selected based on insights from prior research Schaul (2015); Horgan et al. (2018) and preliminary experiments that indicated effective performance within these intervals.

### v2

@@ 1 Introduction
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications Alayrac et al. (2022), Achiam et al. (2023), Zheng et al. (2024), Li et al. (2023). This versatility enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios Yang et al. (2023b), Wang et al. (2024). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
To bridge this gap, fine-tuning(type of light training) aforementioned agents on demonstrations via imitation learning Zhan & Zhang (2023), Nakano et al. (2021), Yang et al. (2023b) channels pre-trained abilities into actionable behaviors suitable for device control. However, the dynamic nature of devices and the web renders models trained on static data sub-optimal as ecosystems evolve. Such agents struggle to recover from mistakes in changing environments, limiting their practical utility. Therefore, building robust and reliable device-control agents necessitates an interactive approach that enables MLLMs to adapt and learn from their own experiences on devices and the Internet.
Moreover, MLLMs are prone to confidently generating incorrect content that deviates from human preferences Ouyang et al. (2022), Ziegler et al. (2019), Bai et al. (2022a), Stiennon et al. (2020). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from Human Feedback (RLHF), becomes essential. RLHF leverages human feedback to align model outputs with desired behaviors and preferences. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to human expectations. Details of the problem formulation as an RL-based fine-tuning approach will be discussed in Sections 2.2 and 3.
@@ 2.1 Multi-Modal On-device Control Agents
Recent advancements in pre-trained Large Language Models (LLMs) and Multimodal LLMs (MLLMs) have revolutionized on-device control agents, moving beyond early methods like behavioral cloning or reinforcement learning Osa et al. (2018), Mnih et al. (2015), Shi et al. (2017). Early agents simulated mouse clicks and typing Shi et al. (2017), Humphreys et al. (2022) but faced scalability and adaptability challenges.
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT Nakano et al. (2021) employ fine-tuned models for web browsing, while WebAgent Gur et al. (2023) generates web code using T5. AppAgent Yang et al. (2023b) and MobileAgent Wang et al. (2024) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding Driess et al. (2023), Reid et al. (2024), but without fine-tuning, they are limited by the base models Driess et al. (2023).
Some works fine-tune VLMs with demonstration data, e.g., AutoUI, CogAgent Kapoor et al. (2024), Zhan & Zhang (2023), but static data-trained models may struggle with real-world variability Jiang et al. (2023). Others use filtered imitation learning with autonomously collected data Pan et al. (2024), Lai et al. (2024), While DigiRL Bai et al. (2024) supports on-device RL fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes. Additionally, DigiRL lacks support for efficient distributed learning algorithm designs, further hindering its scalability and performance in dynamic, parallel settings. To address this, our scalable and asynchronous RL fine-tuning pipeline, DistRL, offers an efficient solution for distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences Stiennon et al. (2020), Ouyang et al. (2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations Ouyang et al. (2022), Bai et al. (2022a).
Recent advances in MLLMs Alayrac et al. (2022), Driess et al. (2023), Reed et al. (2022), Li et al. (2023), Zhang et al. (2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences Ouyang et al. (2022), Ziegler et al. (2019), Bai et al. (2022a), Stiennon et al. (2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative Bai et al. (2022b), Yu et al. (2024), Lee et al. (2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots Bai et al. (2024), Lee et al. (2023), Yu et al. (2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems Reed et al. (2022), Liang et al. (2023), Driess et al. (2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods Driess et al. (2023), Nakano et al. (2021), Reed et al. (2022), Yao et al. (2023). This approach is effective for large-scale applications like Android device control.
@@ 5.1 A-RIDE: The Backbone of DistRL
While GAE has proven highly effective in synchronous environments, its reliance on synchronized data collection makes it less suitable for asynchronous settings, such as distributed control tasks on devices. To handle asynchronous trajectory generation in device control tasks, A-RIDE leverages an enhanced Retrace($\lambda$) method for robust off-policy corrections inspired by Espeholt et al. (2018), Munos et al. (2016). Unlike traditional on-policy algorithms that require synchronous data collection, A-RIDE is designed for distributed environments with asynchronous data.
@@
$\mathcal{L}=-\mathbb{E}_{\mu}\left[\rho_{t}A(s_{t},a_{t})\log\pi(a_{t}|s_{t})\right]-\beta\mathbb{E}_{\mu}\left[\mathbb{H}(\pi(a_{t}|s_{t}))\right]+\lambda\mathbb{E}_{\mu}\left[\mathcal{P}_{\text{invalid}}(a_{t})\right],$ (2)
where the advantage function $A(s_{t},a_{t})$ represents how much better an action $a_{t}$ is compared to the expected reward at state $s_{t}$, given the policy’s understanding of the environment. It can be expressed as the difference between the action-value function $Q(s_{t},a_{t})$ and the state-value function $V(s_{t})$: $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$. The action-value function $Q(s_{t},a_{t})$ estimates the expected return when taking action $a_{t}$ at state $s_{t}$, while the state-value function $V(s_{t})$ estimates the average expected return from state $s_{t}$ under the policy $\pi$. $\mathbb{H}$ is the entropy term, $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini Reid et al. (2024)), ensuring that only contextually appropriate actions are penalized. The hyperparameters $\beta$ and $\lambda$ are optimized through empirical studies to balance exploration and policy robustness effectively. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness.
@@ 5.2 Distributed Prioritized Experience Replay (DPER)
To improve sample efficiency, we employ Distributed Prioritized Experience Replay (DPER). For each trajectory $\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}_{t=0}^{H}$, we compute the priority $p(\tau)$ as: $p(\tau)=w_{1}\overline{|\delta|}+w_{2}\overline{\rho}+w_{3}\overline{\mathbb{H}},$ where $\overline{|\delta|}$ is the average absolute temporal-difference (TD) error over the trajectory, calculated as $\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$; $\overline{\rho}$ is the average importance sampling ratio $\rho_{t}$; and $\overline{\mathbb{H}}$ is the average policy entropy, $\mathbb{H}_{t}=-\log\pi(a_{t}|s_{t})$, encouraging exploration by encouraging policy uncertainty, thus avoiding early convergence to suboptimal policies during training in dynamic environments. The weights $w_{1}$, $w_{2}$, and $w_{3}$ balance the contributions of each component, which is selected by grid-search. Trajectories with higher priorities are replayed more frequently, focusing learning on the most informative experiences. Priorities are periodically updated based on the latest policy, recalculating them to focus learning on the most informative experiences, ensuring continual adaptation to evolving behavior policies. Details can be found in Appendix A.4.
@@ A.4.3 Hyperparameter Tuning for Distributed Prioritized Experience Replay
These ranges were selected based on insights from prior research Schaul (2015), Horgan et al. (2018) and preliminary experiments that indicated effective performance within these intervals.